OpenStudy (anonymous):

Can anybody solve 9x^4+26x^3=3?

OpenStudy (anonymous):

Mmmm.. something's wrong. Here is Wolfram's solution: http://www.wolframalpha.com/input/?i=solve+9x%5E4+%2B+26x%5E3+-3+%3D+0

OpenStudy (anonymous):

How about this? Factorise x^3(9x+26) = 3 3 can only be 3 times 1, or 1 times 3, so x^3 = 3 or 1, and 9x+26 = 1 or 3

OpenStudy (radar):

3 can be 6 times 1/2 as well as many others (including complex values) I suspect there are several values for x

OpenStudy (anonymous):

Attached is the symbolic solution for the two real equation roots. I cannot imagine that this equation would be presented as home work.

OpenStudy (radar):

Geez!, I hope ema123 good luck with it!